Water at 2.5 Mpa and 200°C is heated at constant temperature up to a

quality of 80%. Find (a) the quantity of heat received by the water, (b) the

change in internal energy, and (c) the work of nonflow process. (HINT:

table 4 needed)


1
Expert's answer
2021-11-10T00:49:02-0500

Given T1=200°C, P1=2.5MPa

From the steam tables, at P1=2.5MPa and T1=200°C

u1=849.9 kJ/kg

h1=852.8kJ/kg

At superheated temp,

T2=200°C, ufg=1743.7

x=0.8, hfg=1739.8

u2=u1+xufg

u2=849.9+0.8*1743.7

=2242.42kJ/kg

h2=h1+xhfg

h2=852.8+0.8*1739.8

=2404.1kJ/kg


(a) Q=h2-h1

=2404.1-852.8

=1551.3kJ/kg


(b) ∆u=u2-u1

=2242.42-849.9

=1396.5kJ/kg


(c) Q=∆u+W

From which W=Q-∆u

=1551.3-1386.5

=154.8kJ/kg